OpenStudy (anonymous):

underroot 3x+1- underroot x-1=2

OpenStudy (jiteshmeghwal9):

\[\LARGE{\sqrt{3x+1}}=2+\sqrt{x-1}\]

OpenStudy (anonymous):

and then ?